### v2

# Wisdom of Committee: Distilling from Foundation Model to Specialized Application Model

Abstract
Recent advancements in foundation models have yielded impressive performance across a wide range of tasks. Meanwhile, for specific applications, practitioners have been developing specialized application models. To enjoy the benefits of both kinds of models, one natural path is to transfer the knowledge in foundation models into specialized application models, which are generally more efficient for serving. Techniques from knowledge distillation may be applied here, where the application model learns to mimic the foundation model. However, specialized application models and foundation models have substantial gaps in capacity, employing distinct architectures, using different input features from different modalities, and being optimized on different distributions. These differences in model characteristics lead to significant challenges for distillation methods. In this work, we propose creating a teaching committee comprising both foundation model teachers and complementary teachers. Complementary teachers possess model characteristics akin to the student’s, aiming to bridge the gap between the foundation model and specialized application models for a smoother knowledge transfer. Further, to accommodate the dissimilarity among the teachers in the committee, we introduce DiverseDistill, which allows the student to understand the expertise of each teacher and extract task knowledge. Our evaluations demonstrate that adding complementary teachers enhances student performance. Finally, DiverseDistill consistently outperforms baseline distillation methods, regardless of the teacher choices, resulting in significantly improved student performance.
1 Introduction
The emergence of foundation models, characterized by their immense scale and capacity, has brought remarkable performances Radford et al. (2021); Dosovitskiy et al. (2021); Brown et al. (2020); Devlin et al. (2019); Ramesh et al. (2021). These models, trained on massive amounts of data in an unsupervised fashion, can be adapted to various downstream tasks, either by fine-tuning on application data Hu et al. (2021); Devlin et al. (2019) or leveraging its few-shot or in-context learning ability Brown et al. (2020); Monajatipoor et al. (2023); Bai et al. (2023). Concurrently, efforts in machine learning over the past few years have resulted in specialized application models optimized for specific tasks.
For example, large language models, or LLMs, trained for general language understanding, exhibit impressive performance on recommendation tasks even in the few-shot fashion Kang et al. (2023b). In contrast, recommendation systems based on embedding tables with fully connected layers stand as the default production recommendation system Naumov et al. (2019). LLMs demonstrated strong generalization ability on unseen task data, while recommendation systems based on embedding tables are constructed with domain-specific knowledge and perform well on its optimization distributions. Each category of model has its distinct advantages, and to harness the strengths of both, one natural progression involves transferring knowledge from foundation models into specialized application models, a concept aligned with the goal of knowledge distillation. This process entails the specialized application models learning the task knowledge from the foundation models.
However, transferring knowledge from foundation models to specialized application models presents significant challenges for standard knowledge distillation methods due to the great differences in model characteristics. First, a foundation model often involves an enormous amount of parameters, creating a substantial capacity gap between the teacher and the student. Second, differences in architectures can also be observed. For example, large language models are based on transformer architecture, unlike standard recommendation systems that primarily rely on embedding tables with fully connected layers Kang et al. (2023b); Fan et al. (2023); Naumov et al. (2019). Third, foundation models are typically pre-trained on an extensive corpus of data in a self-supervised manner. In contrast, specialized application models are trained for specific application distributions, leading to a significant distribution gap. Lastly, potential differences in input features and modalities further increase the challenge as foundation models and the student could consume different features. For instance, LLMs can make predictions based on text features such as item descriptions and comments, while standard recommender systems rely on categorical and numerical features such as user history Kang et al. (2023b).
Figure 1: DiverseDistill introduces distillation parameters, specifically, a Question Augmenter and an Answer Augmenter. The Question Augmenter enables the student to understand the teacher and ask tailored questions to each teacher. The Answer Augmenter enables each teacher to reply with a task-oriented answer, which will be used as the soft label in distillation loss.
In this paper, we explore adding complementary teachers who are more similar to the specialized application students in the model characteristics discussed above. A teacher committee with both foundation models and complementary models could bridge the vast differences between foundation models and specialized application models, facilitating a more effective knowledge transfer. Prior research in knowledge distillation also provides supportive evidence; for example, teachers with a smaller gap in model capacity lead to more effective student learning Park et al. (2022); Mirzadeh et al. (2019, 2020). In addition, relying on a single teacher may limit the student’s full potential. This is analogous to forming a Ph.D. committee, which comprises multiple members with diverse expertise, allowing diversity in the teachers to provide unique knowledge and perspective. Foundation models pose a more general understanding, whereas application teachers specialize in downstream tasks.
However, harnessing the benefits of diverse teachers has its challenges.Since the complementary teachers are supposed to be more similar to the student, the teachers (the foundation model teachers and application model teachers) themselves could exhibit considerable differences in capacity, architecture, training distribution, input features, and modalities. For instance, for a rating prediction task where the student employs a standard recommendation architecture, a teaching committee comprising a large language model and a larger recommendation system would teach the student. The recommendation system teacher is more similar to the student regarding architecture, capacity, and data distribution. At the same time, it is evident that these two teachers also possess significant differences. This breaks the common assumption of multi-teacher distillation methods where teachers share similar architecture and training distribution. Under this assumption, standard multi-teacher distillation methods enjoy the performance benefits mainly from a variance reduction perspective, while lack the understanding on teachers’ expertise You et al. (2017); Ding et al. (2022); Du et al. (2020b). Thus, we propose an interactive distillation process, allowing the student to model each teacher’s expertise and ask tailored questions to each teacher. With no constraints on the teachers, DiverseDistill is adaptable to various teacher combinations.
Establishing a teacher committee brings additional distillation costs. Fortunately, the distillation cost from adding a complementary teacher is significantly lower, considering foundation models are highly resource-intensive. Further, with DiverseDistill, each teacher’s importance for the current data is dynamically weighted based on the student’s modeling of the teacher’s expertise. We imagine such weighting mechanics may enable the student to select only one teacher for every training data when distillation cost becomes the major constraint.
In this work, we argue that forming a teacher committee realizes effective knowledge transfer from foundation models to specialized application models. To accommodate a teacher committee without constraints on teacher choices, we design the distillation process as a question-and-answer process with additional parameters for the student to model each teacher’s expertise on the task. Our contributions are summarized as follows.
  • We hypothesize that forming a teacher committee of both foundation models and complementary models would lead to a smoother knowledge transfer from foundation model to specialized application model.
  • In Section  3.2, we introduce DiverseDistill, which learns a distillation module, including a Question Augmenter and an Answer Augmenter, along with the student model. We describe the distillation process in Section 3.1, and how we design and train the distillation module in Section 3.2 and Section 3.3
  • We present our evaluation in Section 4. In Section 4.1, we describe how to enable distillation in recommendation tasks using both the language model and recommendation system as teachers and how to enable distillation in vision tasks using both the vision-language model and vision model as teachers. We demonstrate that leveraging complementary teachers benefits the distillation quality. Also, DiverseDistill significantly increases distillation performance when dealing with a diverse teacher committee, leading to the best student compared to various baselines across different teacher choices.
2 Related Work
2.1 Foundation Model for Specific Application
Foundation models are trained on massive amounts of data and can be applied across various downstream tasks. Example models include BERT Devlin et al. (2019) and GPT Brown et al. (2020) for text, DALLE Ramesh et al. (2021) and CLIP Radford et al. (2021) for image and text, and RT2 Brohan et al. (2023) for robotics. These models can be adapted to downstream tasks via finetuning: updating part of or entire model parameters Church et al. (2021) or adding small additional parameters Hu et al. (2021). Besides, recent foundation models with an immense number of parameters demonstrate impressive zero-shot ability, or in other words, in-context learning ability, which adapts the models by changing the input instead of training the model parameters Dong et al. (2023); Liu et al. (2023).
2.2 Knowledge Distillation
Knowledge distillation is one of the promising directions to reduce the serving cost of deep neural networks Hinton et al. (2015). Knowledge distillation is often more hardware-friendly than orthogonal directions such as pruning Hassibi et al. (1993) and quantization Gholami et al. (2021). However, it has been recognized that significant differences between the teacher and the student create difficulties in knowledge transfer. One of the main differences arises from the capacity in terms of network depth, width, and number of parameters. Recent works in understanding knowledge distillation discover that a large gap between the teacher and the student capacity may lead to degradation in student performance, and one can often employ a teacher of the smaller network to obtain better transferable representation Park et al. (2022); Mirzadeh et al. (2019, 2020). With the bloom of pre-trained models, another major distillation challenge comes from the difference in the tasks. Specifically, when using the pre-trained model as the teacher for downstream tasks, the difference between the pre-training and downstream tasks may lead to a less performant student Zhao et al. (2023); Wu et al. (2023).
2.3 Foundation Model Distillation
The recent bloom of LLMs stimulates an exciting line of works on how to leverage LLM as a teacher to assist smaller language models (Gu et al., 2023; Agarwal et al., 2023; Hsieh et al., 2023). Based on the characteristic in the probability distribution of the auto-regression model, Gu et al. (2023) and Agarwal et al. (2023) design alternative objectives such as reverse Kullback-Leibler divergence. Besides, other works promote LLMs to generate explanations or hard instruction to aid the instruction finetuning process of student models Hsieh et al. (2023).
Before the excitement of LLMs, foundation model distillation focused on obtaining a smaller foundation model with performant transferability. For example, DistillBert Sanh et al. (2020), MT-Bert Wu et al. (2021), and TinyVit Wu et al. (2022) are the distilled version of their original counterpart. The distillation process focuses on maintaining transfer-ability so that the distilled model can be finetuned for different downstream tasks. Researchers also studied distillation from multi-modality foundation models  Sun et al. (2023); Wu et al. (2023); Kim et al. (2023); Wu et al. (2021); Wang et al. (2022a, b). With the student model as a standard multi-modality network, the student mimics both the image and text features from the teacher and improves vision-text benchmarks Wang et al. (2022a, b). Other works in multi-modality distillation focus on the pre-training stage by exploiting the teacher as a pseudo-labeler for unpaired or unlabeled data Sun et al. (2023); Kim et al. (2023).
The above foundation model distillation works illustrated the usefulness of the foundation model in obtaining better students; however, they are explored mainly in a setting with high similarity between the teacher and the student. This paper aims to improve the performance of a specialized application model instead of creating a smaller foundation model. We explore a more general setting in which foundation model teachers and the student model share much fewer characteristics, which opens up more possibilities for using the foundation models.
2.4 Multi-Teacher Distillation
Multi-teacher distillation extends from standard distillation by leveraging multiple teachers to improve student performance under the philosophy of “two heads are better than one". Previous works focus on various ways to combine teachers’ soft targets, logits, or intermediate representations You et al. (2017); Fukuda et al. (2017); Park & Kwak (2019); Asif et al. (2020); Cai et al. (2022) such as averaging over all teachers or randomly selecting one teacher. Further, teacher performances such as task loss or prediction consistency are used as references to assign weights to each teacher Wu et al. (2021); Kang et al. (2023a); Du et al. (2020b); Liu et al. (2020). Moreover, extra information from the student, such as gradient or validation loss, is also used to reflect the usefulness of each teacher and balance the diversity Yuan et al. (2020); Ding et al. (2023); Du et al. (2020a). Besides, methods from reinforcement learning are also applied to score each teacher Yuan et al. (2020).
These methods often assume similar architectures or training distributions among teachers, varying mainly on capacities. In this work, we explore a more challenging setting where teachers can be different regarding capacities, architectures, training distributions, and input features. Such a setting brings more freedom in teacher choices and allows more diversities in teacher expertise.
(a) Question Augmenter
(b) Answer Augmenter
Figure 2: The design of the Distillation Module with the Question Augmenter and the Answer Augmenter. The Question Augmenter introduces Teacher Embedding $E$ to model each teacher’s expertise, and generate questions $\mathbf{q}$ conditioned on $E$. The Answer Augmenter consists of $n$ MLPs, one for each teacher. The Answer Augmenter outputs a set of answers ${a_{i}}$, which will be used as the distillation target.
3 Distilling from Diverse Teachers
We now introduce DiverseDistill, a general distillation framework that learns a distillation module along the student model $\mathcal{S}$ to support distillation from a diverse set of $n$ teacher networks $\{\mathcal{T}_{1},...,\mathcal{T}_{n}\}$. In Section 3.1, we describe how the student learns from the teacher committee with the aid of the distillation module. In Section 3.2 and Section 3.3, we introduce the architecture design and training objective for the distillation module, respectively.
3.1 Learning with the Distillation Module
DiverseDistill uses a two-way communicative process, in contrast to the standard one-way communication in distillation, where the teacher directly provides features or logits as the soft label to the student. With DiverseDistill, the student first asks tailored questions to each teacher, and each teacher then replies to the question with an answer, which will be used as the soft label.
Intuition Continue with the analogy of a Ph.D. committee, as each teacher has different expertise; formulating the question of interest in the language of the expertise area often leads to more accessible communication and more effective advice. Thus, to gain helpful instructions given teachers’ different expertise, DiverseDistill enables each teacher to answer the student’s tailored questions, which are augmented based on the student’s understanding of each teacher.
Distillation Process The distillation module has a Question Augmenter $\mathcal{F_{Q}}$ and an Answer Augmenter $\mathcal{F_{A}}$, as depicted in Figure 1. The process can be divided into three stages.
  1. 1.
    Asking questions. Student model’s hidden states $x$ will be used as the input to the Question Augmenter $\mathcal{F_{Q}}$. The hidden states are expected to appropriately reflect the student model’s current understanding of the data. For example, one such candidate is the [CLS] token in the transformer-based model, which serves as the representation of the entire image/sentence. Then, the Question Augmenter $\mathcal{F_{Q}}$ outputs a set of questions $\mathbf{q}=\{q_{1},...,q_{n}\}$, where one question for each teacher.
  2. 2.
    Collect answers. Each question will be fed into its corresponding teacher model from the early layer, such that each teacher could run a forward pass on its question respectively. Then, the Answer Augmenter $\mathcal{F_{A}}$ takes every teacher’s last layer hidden states and outputs a set of answers $\mathbf{a}=\{a_{1},...,a_{n}\}$.
  3. 3.
    Learning from the answers. The student is trained in the standard distillation fashion, taking the task loss on the original data and the distillation loss on the answers $\mathbf{a}$. To optimize over the distillation dataset $\mathcal{D}$
    $\min\mathbb{E}_{(d,y_{true})\sim\mathcal{D}}(L_{T}(\mathcal{S}(d),y_{true})+\alpha{}L_{D}(\mathbf{a},h))$
    where $L_{T}$ is the task loss, $L_{D}$ is the distillation loss between the answers $\mathbf{a}$ and student’s last layer hidden states $h$, and $\alpha{}$ is the hyper-parameter for loss weight. Inspired by feature distillation, DiverseDistill applies the distillation on the hidden states instead of the logits to handle situations when there are differences in the output space between the teacher and the student. For example, a language model may predict logits over its vocabulary for a recommendation task, while a standard recommendation system predicts logits over the candidate products.
3.2 The Design of the Distillation Module
Question Augmenter DiverseDistill introduce Teacher Embedding Dot Product in the Question Augmenter to generate tailored questions based on each teacher’s expertise. Teacher Embedding Dot Product allows the student to model each teacher’s expertise and generate questions conditioned on the learned teacher embedding.
As shown in Figure 2(a), the Question Augmenter consists of a Teacher Embedding Dot Product and multiple MLPs. The goal of each MLP is to match the model dimension for each teacher, which can be dramatically different with a diverse set of teachers. The Teacher Embedding Dot Product takes the student model’s representation of data $x$ as input and yields a set of questions with hidden states. The following MLPs outputs questions $\mathbf{q}=\{q_{1},...,q_{n}\}$ based on the questions hidden states.
[…]
Teacher Embedding Dot Product also provides a way to dynamically weight each teacher based on the teacher embedding and student data representation. Specifically, we compute the teacher importance $w\in\mathbb{R}^{n}$ as $w=\sigma(FE^{T})$, where $\sigma$ is the Sigmoid function and $F=qW_{f}$.
Answer Augmenter The Answer Augmenter takes a set of hidden states $\mathbf{p}=\{p_{1},...,p_{n}\}$, one from each teacher as the input, as depicted in Figure 2(b). Each hidden state will be passed through a separate MLP to match back with the student model dimension. The Answer Augmenter yields a set of answers $\mathbf{a}=\{a_{1},...,a_{n}\}$, which will be used as the soft label in distillation.
Distill from the Answers DiverseDistill computes the distillation loss $L_{D}$ between the student model’s last hidden states $h$ with the set of answers $\mathbf{a}$, dynamically weighted by the teacher importance $w$. Formally,
$L_{D}(\mathbf{a},h)=\sum_{1}^{n}w_{i}L(a_{i},h)$
where $L$ denotes a divergence metric such as $L_{2}$ or the Kullback-Leibler divergence. We use $L_{2}$ in our evaluation.
3.3 Training the Distillation Module
DiverseDistill learns the Question Augmenter and the Answer Augmenter with the student. We introduce a task regularizer to encourage the distillation module to generate task-oriented questions and answers.
Challenges The challenges arise as we make no assumptions on the teacher committee. With a diverse set of teachers, the hidden states on the same input can be dramatically different. When asking questions, as the teacher model needs to run inference pass on the questions, questions must be aligned with each teacher and appropriately query the teacher’s knowledge on the task. When learning from the answers, the information embedded in teachers’ hidden states can be excessive and biased for the student model, especially when considering teachers with capacity, architecture, and distribution gaps.
Figure 3: Illustration for the task regularizer. DiverseDistill takes losses on the teacher’s predictions based on the questions and the student’s predictions based on the answers.
Task Regularizer We introduce task loss on the predictions based on the questions and answers to align different models’ hidden states and regulate the distillation process to be task-oriented. The intuition is that the question has to be aligned with the teacher model to the extent that the teacher can make the correct prediction. Similarly, the answer should be digestible bites such that the student can make the correct prediction based on the answer.
We visualize the process in Figure 3 with one teacher as an example. The hidden states after the student model’s early layers are fed into the Question Augmenter. The teacher model takes the question as the input for the intermediate layers. The hidden states after the teacher’s intermediate layers are fed into the Answer Augmenter; then, the student takes the answer as the input for the last layer. This process yields two predictions: $\hat{y}_{Q}$ is the teacher prediction using tailored questions, and $\hat{y}_{A}$ is the student prediction with injected answers. The parameters in the Question Augmenter and the Answer Augmenter are optimized for
$\min\mathbb{E}_{x\sim\mathcal{X}}(L_{T}(\hat{y}_{Q},y_{true})+L_{T}(\hat{y}_{A},y_{true}))$
where $L_{T}$ is the task loss function. Note that the gradient from the task regularizer is only applied for the distillation module, as we freeze the student and teacher models and mark them in grey in Figure 3.
4 Experiment
We evaluate DiverseDistill over a diverse set of datasets covering both recommendation and vision applications. We describe the experimental setup in Section 4.1 and baselines in Section 4.2. In Section 4.3, we introduce implementation details. We provide the main evaluation result in Section 4.4. Finally, we investigate the influence from different teacher committees and other design choices in Section 4
4.1 Experiment Setup
The goal of this evaluation is to answer two questions:
  • Do complementary models benefit the distillation process from foundation models?
  • Is DiverseDistill effective in accommodating a diverse set of teachers?
Thus, for both recommendation and vision tasks, we choose one foundation model as teacher and other models that are more similar to the student model in model architecture, training data, and input feature as complementary teachers. We describe the committee choices separately for each application:
  • Rating Prediction. We evaluate on standard benchmark datasets MovieLens Harper & Konstan (2015). The student model is an MLP model with embedding tables, following standard recommendation architecture Naumov et al. (2019).
    The foundation model teacher is a large language model. Recent works have shown promising results of large language models in recommendation tasks Fan et al. (2023); Kang et al. (2023b). We follow Kang et al. (2023b) to formalize the input features in natural language and finetune T5 model Raffel et al. (2023) on the task dataset. For example, input is presented as “user u-138, a doctor. Watched movie ‘One Flew Over the Cuckoo’s Nest (1975)’, id m-357. Rating: 5".
    We want complementary models to be more similar to the student model. Thus, we include MLP-based recommendation models of larger capacity and train them from scratch on the task dataset. Note that the MLP-based recommendation model and language model use different features here. MLP-based recommendation models use embedding tables and employ categorical features such as user ID and movie ID. On the other hand, large language models operates with natural languages and utilize text features such as movie titles. We summarize the teacher details on size, architecture, and input features in Table 1.
  • Image Classification. We evaluate on standard image benchmark datasets CIFAR-10, CIFAR-100 Krizhevsky (2009), and ImageNet Russakovsky et al. (2015). The student model is a Vision Transformer (ViT) Dosovitskiy et al. (2021).
    The foundation model teacher is a pre-trained text-vision model, CLIP Radford et al. (2021). Following the original paper, we use CLIP for image classification under the zero-shot setting. The complementary model is a larger Vision Transformer (ViT) finetuned on the task datasets. Note that CLIP uses both the text of the class label and the image as input, while the input for the student is only the image. The complementary teacher is closer to the student in model architecture, training distribution, and input features.
Table 1: Details on the student and teachers regarding architecture, capacity, and input features for recommendation tasks.
Model Size Architecture Input Features
Student MLP-S 2M Embedding Table + MLP [128, 64] Categorical
Teacher T5 76M Encoder-Decoder Language Model Categorical + Text Feature
MLP-M 3M Embedding Table + MLP [512, 256] Categorical
MLP-L 4M Embedding Table + MLP [1024, 512] Categorical
Table 2: Details on the student and teachers regarding architecture, capacity, and input features for vision tasks.
Model Size Input Features
Student Vision Transformer (ViT) 86M Image
Teacher CLIP 149M Image + Text
Vision Transformer (ViT) 307M Image
4.2 Baselines
We consider state-of-the-art distillation methods from foundation model distillation and multi-teacher distillation to answer the above two evaluation questions.
From foundation model distillation, we organize the baselines in terms of methodologies as some works share the same methodology but were proposed for different foundation models or different applications with slight adjustments to accommodate the architectures.
  • Logit Distillation (LD): The logits of the teacher model are used as the soft target. This is widely used in different foundation model distillation works Sanh et al. (2020); Wu et al. (2022).
  • Feature Distillation (FD): The hidden states of the teacher model are used as the soft target. Projection layers are added to bridge the dimensional differences between the teacher and the student. This is widely used in different foundation model distillation, especially from multi-modality foundation models Sun et al. (2023); Wang et al. (2022b).
  • Student Customized Knowledge Distillation (SCKD) Zhu & Wang (2021): SCKD examines the alignment between gradient from distillation loss and that from task loss.
From multi-teacher distillation, we consider the following:
  • MultiTeacher (MT) You et al. (2017): MT simultaneously incorporates multiple teacher networks by averaging the soft labels.
  • Adaptive Multi-Teacher Multi-Level Knowledge Distillation (AMTML) Liu et al. (2020): AMTML learns from the intermediate layers of each teacher using a multi-group hint strategy.
  • Stochastic Knowledge Distillation to obtain compact BERT style language model (AutoSKDBert) Ding et al. (2022): AutoSKDBert learns a weight for each teacher based on student progress.
  • Adaptive Ensemble Knowledge Distillation (AEKD) Du et al. (2020a). AEKD dynamically weights teacher models depending on the agreement in the gradient space.
Table 3: This table summarized the student performance improvement given different single teachers and distillation methods. The percentage of performance improvement is measured against the best teacher performance without distillation. Underline shows the best performance of all BASELINE methods, while bold is the best performance across all methods.
MovieLens CIFAR-10 CIFAR-100 ImageNet
T5 MLP CLIP ViT CLIP ViT CLIP ViT
LD NA 54.4% -15% 30% 40.8% 42.3% 24.7% 20.8%
FD 31.4% 18.2% 20.0% 32.5% 42.0% 34.3% 22.1% 20.8%
SCKD 28.2% 53.8% 28.8% 31.3% 7.5% 10.4% 25.3% 22.7%
DiverseDistill 38.9% 73.1% 37.5% 36.3% 49.8% 42.8% 79.9% 81.2%
Table 4: This table summarized the student performance improvement given a teacher committee with different multi-teacher distillation methods. Underline shows the best performance of all BASELINE methods, while bold is the best performance across all methods.
MovieLens CIFAR-10 CIFAR-100 ImageNet
T5 + MLP CLIP+ViT CLIP + ViT CLIP+ViT
MT 60.18% -12.5% 43.3% 21.4%
AMTML 30.73% 13.7% 29.4% 22.1%
AutoSKDBert 63.45% -10.0% 47.8% 77.9%
AEKD 43.27% 31.3% 15.4% 24.7%
DiverseDistill 114.18% 82.5% 73.1% 90.9%
4.3 Implementation Detail
In this section, we discuss in details for each application (1) what is used as the input to the Question Augmenter, (2) what is used as the input to the Answer Augmenter, and (3) how do each teacher process the questions.
  • Recommendation tasks. The student model is a standard recommendation system with embedding tables and MLPs. We use the representations retrieved from the embedding table as the input to the Question Augmenter. We use the hidden states before the last layer in teacher models as the input to the Answer Augmenter.
    Teacher models include both standard recommendation systems and language models. For recommendation system teachers, we replace the hidden states after the embedding table with the generated questions. For language teachers, we replace the token embedding after the first transformer layer in the decoder with the generated question at the decoding stage.
  • Vision tasks. The student model is a Vision Transformer, and we use the representations of the classification token from the middle layer as the input to the Question Generator.
    Teacher models include both CLIP and Vision Transformer. For the Vision Transformer teacher, we use the representation of the classification token from the last layer as the input to the Answer Augmenter. For CLIP teachers, we utilize the representation from the image encoder as the input to the Answer Augmenter. We replace the classification token in the middle transformer layer for the Vision Transformer teacher with the generated question. We replace the classification token in the transformer image module for the CLIP model with the generated question.
4.4 Main Results
We conclude our main experiment results in Table 3 and Table 4. For easier representation given different tasks and teachers, we report the percentage of performance improvement against the best teacher performance without distillation. For example, in MovieLens, single MLP teacher performs better than T5 teacher. Thus, we measure the performance improvement of different distillation methods in Table 3 against the performance of MLP (without distillation) for MovieLens. Raw task evaluation number can be found in Appendix A
Comparing Teacher Choices In Table 3, we summarize the teaching performance from the foundation model compared to other teacher choices on a single-teacher basis. We have the following observations: (1) with different standard distillation methods, i.e. baselines, the foundation model teachers fail to transfer their own superior performances to the students. In contrast to the high expectations on foundation models, we observe that on most datasets, smaller teachers trained on the task data lead to better students, as marked with underlining. The only exception is ImageNet, where CLIP and ViT teachers perform similarly. This is because CLIP is pretrained on ImageNet, which doesn’t exhibit as much distribution gap as other datasets. (2) even for single teacher distillation, DiverseDistill can still boost student performance, regardless of the teacher’s choices. Specifically, we observe that on CIFAR-10 and CIFAR-100 datasets, DiverseDistill from CLIP generates the best student performance.
Accommodating Diverse Committee In Table 4, with complementary teachers, we observe more significant variances in the distillation performances with multi-teacher baselines, due to the diversity in the teacher committee. Comparing numbers with underlines, i.e., the best performer among baselines, between Table 3 and Table 4, we find that even without applying DiverseDistill, having a teacher committee generally leads to a better student. When applying DiverseDistill, we observe consistent and significant improvements compared to all multi-teacher baselines across all datasets.
Adding the two pieces together, using DiverseDistill for both foundation models and complementary models as teachers significantly improves teaching performance, leading to the best student in all scenarios.
4.5 Ablation Study
In this section, we present additional results investigating different components of this paper.
Similar Committee vs. Diverse Committee We compare the a committee of diverse teachers against a committee of similar teachers. Specifically, we train two MLP recommendation systems of different sizes to form a teacher committee and compare with a teaching committee of a foundation model plus an MLP recommendation system. As shown in Table 5, we observe a better student with T5 and MLP as teachers. This confirms the benefit of both the foundation model and diversity.
Table 5: This table summarizes the student performance improvements given two different teacher committees, a more diverse one (T5 + MLP) vs. a more similar one (two MLPs).
MovieLens
MLP-M + MLP-L T5 + MLP-L
DiverseDistill 93.27% 114.18%
Figure 4: Visualization of the importance score of 11 random data samples from MovieLens.
Selecting Teacher based on Importance Score While we showcase significant improvements in distillation quality, one drawback involves the increasing computation cost. One potential approach to mitigate the increased cost due to the growing number of teachers is to select the most suitable teacher for each data based on the importance score.
The importance score is designed to reflect the usefulness of each teacher for every data sample. One can use the importance score for teacher selection to reduce the distillation cost resulting from multiple teachers. Specifically, students only ask questions to the teacher with a score greater than a threshold, which saves the forward computation cost of the teachers who are not selected. We form a teacher committee with T5 and two MLP recommenders (i.e. MLP-L and MLP-M) and visualize the importance score of 11 random data points in Figure 4. When using 0.6 as the threshold, we observe no loss in the distillation performance while saving around 30% of teachers’ passes.
5 Conclusion
One remarkable aspect of foundation models is their ability to excel across different tasks. And to facilitate intelligence on various tasks under a serving cost, it is crucial to transfer the task knowledge from a giant foundation model to a more efficient task model. In pursuit of this goal, we study distillation from foundation models where differences in architecture, data distribution, modality, and capacity exist between the student and teacher, mirroring the substantial gap between the foundation model and the downstream task model. Our contribution unfolds in two ways. First, we demonstrate that enhancing distillation performance is possible by incorporating complementary teachers and forming a teaching committee. Second, we introduce DiverseDistill, which is a general and promising distillation method applicable for distillation from a diverse set of teachers.
Appendix A Additional Experiment Results
In Section 4, we report performance improvement, which is more straightforward to reflect distillation quality, especially given both recommendation and vision tasks. Here, we provide model accuracy for reference. Specifically, we report Mean Square Error on MovieLens, and classification accuracy on CIFAR-10, CIFAR-100, and ImageNet.
Table 6: This table summarized the student accuracy given different single teachers and distillation methods. Mean Square Error is reported for MovieLens, and classification accuracy is reported for CIFAR-10, CIFAR-100, and ImageNet.
MovieLens CIFAR-10 CIFAR-100 ImageNet
T5 MLP CLIP ViT CLIP ViT CLIP ViT
LD NA 0.8461 0.9759 0.9795 0.9223 0.9226 0.7758 0.7752
FD 0.8587 0.866 0.9787 0.9797 0.9226 0.921 0.7754 0.7752
SCKD 0.8605 0.8464 0.9794 0.9796 0.9156 0.9162 0.7760 0.7755
DiverseDistill 0.8546 0.8358 0.981 0.98 0.9241 0.9227 0.7840 0.7845
Table 7: This table summarized the student accuracy given a diverse teacher committee with different multi-teacher distillation methods. Mean Square Error is reported for MovieLens, and classification accuracy is reported for CIFAR-10, CIFAR-100, and ImageNet.
MovieLens CIFAR-10 CIFAR-100 ImageNet
T5 + MLP CLIP+ViT CLIP + ViT CLIP+ViT
MT 0.8429 0.9761 0.9228 0.7753
AMTML 0.8591 0.9782 0.920 0.7754
AutoSKDBert 0.8411 0.9763 0.9237 0.784
AEKD 0.8522 0.9796 0.9172 0.7758
DiverseDistill 0.8132 0.9837 0.9288 0.787